Bramble Makeup produces face cream. Each bottle of face cream costs $10.50 to produce and can be sold for $12.90. The bottles can be sold as is, or processed further into sunscreen at a cost of $14.60 each. Bramble Makeup could sell the sunscreen bottles

Bramble Makeup produces face cream. Each bottle of face cream costs $10.50 to produce and can be sold for $12.90. The bottles can be sold as is, or processed further into sunscreen at a cost of $14.60each. Bramble Makeup could sell the sunscreen bottles for $22.40 each. What should Bramble Makeup do?

a. Face cream must be further processed because its profit is $7.80 each.

b. Face cream must not be further processed because costs increase more than revenue.

c. Face cream must not be further processed because it decreases profit by $2.70 each.

d. Face cream must be further processed because it increases profit by $2.40 each.
Answer #1

ans. B 

face cream revenue= 12.90 - 10.50 = 2.40

sunscreen revenue = (10.50+14.60) - 22.40 



answered by: Raghuinderbir Singh Mann